For the given state of plane stress, calculate the strains on each faces. (E=200 GPa, G=70 GPa, and v=0.3). a) Calculate the principal strains and angle ?? at which it occurs. b) Calculate the maximal shear stress in the plane and angle ?? at which it occurs. (Hint; You can construct Mohr’s Circle for strain transformation with the same rules as we constructed in stress transformation)80 MPa 40 MPa 50 MPa
Answer #1

DATE tan (20p) = 2xy = 2x 40 6x-by - 50-80 2x60 tan (20p) - 0.615 -130 20p = 31.60 Qp = 15.800 Os = 15.80 + 45° = 60 and 29.2Please go through the answer